OpenStudy (anonymous):

If f(x) = -x - 2 and g(x) = x2, what is (g º f)(-6)? HELP PLEASE?!

OpenStudy (studygurl14):

g of f (-6) = g(f(-6)) Solve for f(-6) first, then take the answer from that, and plug it in for x in g(x)

OpenStudy (anonymous):

what does f equal and g ?? im confused

OpenStudy (studygurl14):

Look here: \[f(x) = -x - 2\] \[g(x) = x^2\] \[(g ºf)(-6) = g(f(-6))\] \[f(-6) = -(-6) -2 \] Solve and whatever value you get for f(-6), plug in for x in the equation of g(x), and solve. That's your answer

OpenStudy (anonymous):

thank you for helping. i dont understand it still

OpenStudy (studygurl14):

Maybe this will help: http://www.mathsisfun.com/sets/functions-composition.html